OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2, 8, -26, -80, … -2, 8, -26, 80, … -2, 12, -44, 180, … -2, -12, -44, -180, …

OpenStudy (kc_kennylau):

Don't be lazy :p, find f(3) recursively :)

OpenStudy (kc_kennylau):

lol ok i find it for you

OpenStudy (kc_kennylau):

\(f(3)=-4f(2)-3f(1)=-4[-4f(1)-3f(0)]-3f(1)=-4[-4(8)-3(-2)]-3(8)=80\)

OpenStudy (anonymous):

Thank you so much

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (kc_kennylau):

lol didn't realize that it was over the margin sorry I actually wrote \(=80\) at the end but it won't show it in my browser :/